OpenStudy (anonymous):

Mr. Hauseman has 17 students in his class, three of which are freshman and the others are sophomores and juniors. He is going to randomly draw two students to be partners. He calculates the probability of drawing first a freshman and THEN a sophomore to be 3 34 . How many sophomores are in the class?

OpenStudy (kropot72):

Let x be the number of sophomores. The probability of drawing first a freshman is 3/17. The probability of drawing a sophomore next is x/16. Therefore we can form the equation: \[\frac{3}{17} \times \frac{x}{16}=\frac{3}{34}\ ..........(1)\] Now all you need to do is solve equation (1) to find x.